The average of six daily high temperatures is 58˚ Fahrenheit. If two of the temperatures were 48˚ and 32˚, what was the average of the other four temperatures?

Answer by mananth(16949)

the average of 6 high temperatures = 58
the sum of these temperatures = 6 * 58= 348
..
the sum of temperatures on other two days = 48+32=80
348-80 = sum of temperatures on the other 4 days.
=268
Average for 4 days = 268/4
=47 deg